For the diprotic weak acid H2A, K a1 = 3.1 × 10 − 6 and K a2 = 7.2 × 10 − 9 . What is the pH of a 0.0500 M solution of H 2 A ?

pH =

What are the equilibrium concentrations of H 2 A and A 2 − in this solution?

[ H 2 A ] =

[ A 2 − ] =
Answer #1


To find the pH of a 0.0500M solution of HLA Given that Ku = 3.1x10 Kg = 7.2x104 It will dissociates HA HA +H* K1 = 3.1x10 HAH

To find the[HA] [HA]=0.0500- x [HA]=0.0500 -0.000393 = 0.0496M To find the [A-] For the secondreaction HA A A2 +H* Ka = 7.2x1